The position of a particle moving along the x-axis is determined by the

equation

d^2x/dt^2+8x=20cos2t

. If the particle starts from rest at x = 0,


determine

(i) The position x as a function of time


1
Expert's answer
2020-11-16T07:49:24-0500

Solution

This is second order differential equation

"\frac{d^2x}{dt^2}+8x=20\cos 2t"

Firstly found it's

CF and PI

"x(t) =CF+PI"

CF is

(m2+8) =0

it's solution become

"m=-2\sqrt{2}i \\m=+2\sqrt{2}i"

So CF="(A\cos2\sqrt{2}t+B\sin2\sqrt{2}t)"

And

PI="\frac{20\cos2t}{(D^2+8)}=\frac{20\cos2t}{(-4+8)}=5\cos2t"

Therefore

Position become

"x(t) =(A\cos2\sqrt{2}t+B\sin2\sqrt{2}t)+5\cos 2t"